OpenStudy (anonymous):

The ratio of the areas of two similar polygons is 225:25. What is the ratio of their corresponding sides? 2:1 3:1 5:1 9:1

OpenStudy (anonymous):

LOL that is not you in that picture

OpenStudy (nurali):

Ratio of areas = a^2: b^2 Ratio of sides = a:b In this case... Ratio of areas = (15)^2: (5)^2 Ratio of sides = 15:5 Ratio of sides = 3:1

OpenStudy (anonymous):

Ok, I thought 3:1 as well, I will let you know..

OpenStudy (anonymous):

Can you help me with one more tough question?

OpenStudy (anonymous):

Two regular pentagons have areas of 20 mm² and x² - x mm². What values of x make the pentagons congruent? 4 and 5 4 and -5 -4 and 5 -4 and -5